A manufacturer of golf equipment wishes to estimate the number of left-handed golfers. How large a sample is needed in order to be 98% confident that the sample proportion will not differ from the true proportion by more than 5%? A previous study indicates that the proportion of left-handed golfers is 8%.
Answer #1

Solution :

Given that,

\hat p = 8% = 0.08

1 - \hat p = 1 - 0.08 = 0.92

Margin of error = E = 5% = 0.05

At 98% confidence level the z is ,

\alpha = 1 - 98% = 1 - 0.98 = 0.02

\alpha / 2 = 0.02 / 2 = 0.01

Z\alpha/2 = Z0.01 = 2.326

Sample size = (  Z\alpha/2 / E)2 * \hat p * (1 - \hat p)

= (2.326 / 0.05)2 * 0.08 * 0.92

= 159.27 = 160

Sample size = n = 160